# Toward Green and Human-Like Artificial Intelligence: A Complete Survey on Contemporary Few-Shot Learning Approaches

### v2

@@ 3.1.4. Probabilistic Extensions
To alleviate deterministic meta-learning’s inability to quantify uncertainty, various probabilistic approaches have been developed to improve performance and uncertainty quantification by modeling parameters as random variables and obtaining their distributions. In this direction, Bayesian meta-learning is a popular paradigm that fits expressive generative models to small datasets by incorporating inductive priors learned from related tasks. One of the first works to explore probabilistic FSL is (Edwards and Storkey, 2017), where a hierarchical version of Variational Autoencoders (VAEs) is proposed that can be used for FSL by learning the statistics of the training data and then using them to classify new samples. The basic idea is to perform variational inference at two levels using a double variational bound where a higher-level latent variable describes datasets and a lower-level latent variable describes individual examples.
@@
Additionally, many subsequent works have focused on extending MAML to the probabilistic setting. In (Grant et al., 2018), MAML is reinterpreted as an inference for the parameters of a prior distribution in a hierarchical Bayesian model, and it is shown that the choice of update rule in the inner loop corresponds to a choice of prior over task-specific parameters. To mitigate MAML’s inaccurate point approximation, a Bayesian variant of MAML using the Laplace approximation is also proposed. In (Finn et al., 2018), MAML is extended by learning the joint posterior of initial and task-specific parameters, conditioned on each task’s support set, using structured variational inference. An amortized variational inference technique is also proposed to avoid storing separate variational distributions for each task’s initial and task-specific parameters. A Bayesian variant of MAML leveraging non-parametric variational inference has also been proposed in (Yoon et al., 2018), which learns the posterior distributions of task-specific parameters using Stein Variational Gradient Descent to approximate the posterior over the weights in the final layer with a particle ensemble. However, the posterior’s expressiveness depends on the number of particles in the ensemble, leading to poor scaling regarding memory and computation costs.
@@
Apart from probabilistic MAML extensions, novel architectures have also been proposed, such as deep kernels where both feature extractor and base kernel parameters are meta-learned across tasks(Patacchiola et al., 2020). The deep kernel’s parameters are learned in the outer loop by maximizing the marginal likelihood across all tasks using a Gaussian process approach, obviating the need for an inner loop and leading to an efficient and straightforward closed-form solution. Finally, a novel method is introduced in (Pavasovic et al., 2023) that meta-learns a Bayesian prior on the function space to avoid overconfident estimations. The problem is formulated as a stochastic process with different realizations, and the prior’s score function is estimated using a transformer encoder architecture.
@@
Lastly, probabilistic meta-learning can also mitigate some of metric-based meta-learning’s limitations. Specifically, metric-based meta-learning methods typically obtain a single point estimate for each class, leading to biased estimates due to a lack of data. However, a proposed extension in (Zhang et al., 2019) models each class with its distribution through variational inference, allowing for the calculation of confidence and more accurate classification of novel samples. Expanding on this, in (Sun et al., 2021), computational complexity is reduced by using fewer random variables for latent class prototypes. A global model is used to learn the posterior distributions of these variables through amortized inference, and task-specific estimates of variational parameters are obtained using initial parameters learned across tasks.
3.2. Transfer Learning Approaches
@@
Figure 6. Two-stage model finetuning in transfer learning.
@@
One of the first works to explore the efficiency of finetuning techniques for FSL has been in (Chen et al., 2019), where it is argued that finetuning methods’ performance has been underestimated in the past. Consequently, a simple baseline that utilizes a pretrained feature extractor backbone and finetunes only the classification head within each few-shot task is proposed. Classification is then performed based on the cosine similarity between the prototypes extracted from the weight matrix of the classifier head and the feature representations of the given samples. Similarly, in SimpleShot(Wang et al., 2019), a pretrained frozen backbone is also used to extract feature representations. However, a series of simple transformations such as centering and $\ell_{2}$-normalization are applied to the extracted representations, and the final classification is performed using the nearest neighbor/centroid algorithm.
@@
On a different approach, in (Wang et al., 2022b), it is suggested that transferring knowledge from a base dataset to new tasks without retraining the pretrained backbone could lead to biased results due to the extraction of distorted features for out-of-distribution samples. To avoid this and prevent overfitting of the feature extractor due to the small number of samples, the whole model is finetuned using an inverted Firth Bias regularizer that regularizes the distribution of the predicted labels to be more uniform, similar to label smoothing. Full model finetuning is also proposed in (Dhillon et al., 2020). In particular, the pretrained backbone is combined with a linear classifier head, and a transductive objective is introduced to finetune both the backbone and the classifier within each task, granting it the ability to adapt to different numbers of ways and shots without retraining.
@@
To address both limitations of feature distortion using a frozen pretrained backbone and the possibility of overfitting when finetuning the whole model, partial finetuning, e.g., finetuning only specific layers of the backbone network, has been proposed in (Shen et al., 2021). To determine the finetuned layers, a genetic programming algorithm is used to learn each layer’s learning rates during finetuning, automating the whole procedure. On the other hand, in (Shysheya et al., 2023), FiLM modules are introduced to modulate the backbone network without retraining all of its parameters, substantially reducing the number of updated parameters. On top of it, a Naive Bayes classifier is used for prediction.
@@
Finally, various other approaches have also emerged, aiming to tackle FSL via model finetuning. In (Rizve et al., 2021), the effect of invariance and equivariance is explored in the context of learning better feature representation of samples for FSL. Subsequently, the authors propose learning two additional modules alongside the classification head for enforcing equivariance and invariance via self-supervised learning, and the final model is obtained via self-distillation. Transfer learning has also been combined with meta-learning to bridge the gap between the two methods. Specifically, n (Demirel et al., 2023), meta-learning is used to tune an object detection model’s hyperparameters, while finetuning is used to tackle the few-shot object detection objective in each task.
@@
Overall, finetuning methods can be a great asset in few-shot applications where simplicity and efficiency are crucial. However, there is no clear consensus on whether using a frozen backbone or finetuning it either fully or partially leads to better performance(Luo et al., 2023). Additionally, pretraining the backbone network in a large base dataset is not always possible in some domains. Finally, cautiousness is needed when there is a domain gap between the base dataset and the novel tasks since it could hinder learning due to negative transfer.
3.3. Data Augmentation Approaches
@@
3.3.1. Basic Data Augmentation
@@
The main logic behind these approaches is to regularize the model by preventing it from overfitting to particular regions of an image and ignoring other parts. In general, this is achieved by removing specific parts of an image and replacing them with new ones, forcing the model to focus on different features. For instance, Cutout randomly removes square regions (cutouts) of the training images, while Random Erasing randomly selects and erases rectangular regions within an image and replaces them with random or predefined pixel values. On the other hand, a more recent approach, Context-Guided Augmentation(Chen et al., 2023a), adds new instances of classes strategically using contextual information to simulate occlusion occurrences, with each segmented instance being selectively positioned onto the object’s bounding box in a randomized manner. Overall, image erasing methods can be seen as a form of spatial dropout that induces robustness against object occlusions and different spatial configurations. However, they suffer from the same limitations as image manipulation methods since they must be hand-picked beforehand.
3.3.2. Advanced Data Augmentation
@@
These methods aim to enhance generalization performance across diverse datasets and mitigate the risk of overfitting by integrating information from multiple sources. In Mixup(Zhang et al., 2018), this is achieved by blending two or more images by taking a convex combination of their pixel values and their corresponding one-hot label encodings to create novel hybrid samples. Notably, these linear interpolations introduce diversity to the training data and assist models in learning more robust and generalized features with minimal computational overhead. CutMix(Yun et al., 2019), on the other hand, extends this concept by cutting and pasting rectangular patches from different images, fostering spatially coherent variations. More recently, Multi-stage Augmented Mixup(Liang et al., 2023) (MiaMix), a novel mixing ratio sampling method, has been proposed, consisting of four stages: random sample pairing, sampling of mixing methods and ratios, mixing mask generation and augmentation, and obtaining the mixed output. MiaMix selects a mixing method from a pool of methods and sets the mixing ratio for each mask. The masks are then augmented and mixed in the final stage. Finally, PatchMix(Xu et al., 2022b) has been proposed to address the existence of biases, such as spurious correlations, in FSL. Specifically, it disentangles causal and non-causal features to improve model invariance and generalizability by augmenting original query images with images from another query while preserving the label of the additional query.
@@
Although hand-crafted DA methods have been invaluable in dealing with data scarcity and creating models invariant to various transformations, manually selecting the optimal augmentations to apply is a challenging task that often leads to suboptimal results. As a result, novel approaches have been suggested to automate this selection process by systematically finding the optimal selection policy. AutoAugment(Cubuk et al., 2019) formulates a search space within which each policy consists of multiple sub-policies that encapsulate different image transformations. In this setting, the most effective policy is identified through a search algorithm that maximizes validation accuracy on a target dataset based on RL principles. On the other hand, Randaugment(Cubuk et al., 2020) automates this selection by introducing a randomized approach. In particular, it dynamically selects a set of augmentation operations from a transformations pool and applies them with random magnitudes to each image. This approach strikes a balance between variability and consistency while it also obviates using an RL algorithm to find the optimal policy.
Figure 7. Feature data augmentation techniques extract the image’s feature representation and apply the transformation in the feature space.
@@
While traditional DA techniques have focused on applying transformations on the raw input data, feature augmentation techniques have also been used to enhance the learning capabilities of a model by manipulating the extracted feature representations. This can involve methods such as introducing new features, transforming existing ones, or modifying specific feature values. For instance, the work in (Wang et al., 2023a) proposes the Information Fusion Rectification method to identify relevant base class prototype features from the query set using cosine similarity. The method leverages the relationship between base, support, and query samples and rectifies the support set’s distribution using cosine similarity metrics as weights, generating new samples on the extracted feature space relevant to the provided task. Figure 7 shows an example of applying feature data transformations.
@@
A different way to perform DA could be by leveraging modern techniques such as Neural Style Transfer, which offers a unique approach focusing on style-based variations. In (Jackson et al., 2019), the use of random style transfer is proposed by introducing randomness to texture, contrast, and color while preserving the underlying shape and semantic content of data. This is achieved by adapting an arbitrary style transfer network to perform style randomization, with target style embeddings being sampled from a multivariate normal distribution instead of being computed from a specific style image. Finally, diffusion models have also been utilized in this context. Specifically, in (Feng et al., 2023), a pretrained CLIP encoder is used to extract latent features fed to Stable Diffusion to generate synthetic images. Data diversity is ensured by filtering generated images based on cosine similarity, while augmentation is enabled in image and feature spaces.
@@
3.4. Neural Processes
Parallel to the development of probabilistic models that extend the classic meta-learning framework, significant research effort has revolved around Neural Processes (NPs), another family of probabilistic models suitable for FSL. On a higher level, NPs aim to bridge the gap between deep neural networks (DNNs) and Gaussian Processes (GPs). DNNs excel at function approximation, yet they require retraining for each new function. On the other hand, GPs can quickly infer new functions at test time, but they are computationally expensive and sensitive to the selection of priors(Garnelo et al., 2018a). NPs can make accurate predictions given a few training data points and scale to complex functions and large datasets. NPs were initially developed as function estimators, suitable for modeling stochastic processes. However, their ability to model different functions using only a few samples makes them ideal for few-shot regression, where each function is a different task, and the context data constitute the support set.
The first proposed model of this family has been Conditional Neural Processes(Garnelo et al., 2018a) (CNPs) that uses a permutation-invariant neural network to parametrize its dependence on the observations. For each task, an aggregated mean representation of the support set is obtained and used by a decoder that models the distribution parameters of the examined target point. The model is trained end-to-end through gradient descent to maximize the conditional likelihood of a random subset of targets given a random observation set. Extending CNPs, NPs(Garnelo et al., 2018b) introduce a more generalized approach by using a global latent variable that allows for better uncertainty modeling. Similar to CNPs, an aggregated mean representation of the support set data is obtained, which conditions the global latent variable. Finally, the decoder receives the target data and a sampled value from the conditioned global latent variable as input and predicts the output values.
Yet, despite their desired properties, these methods lack expressivity power and suffer from underfitting in complex settings. To address these limitations, Attentive Neural Processes(Kim et al., 2019) (ANPs) introduce an attention mechanism to enhance model performance. Specifically, self-attention is used to capture the interactions between support set samples, and cross-attention is used to obtain the aggregated representation. However, these attention modules also lead to increased computational complexity. Convolutional Conditional Neural Processes(Gordon et al., 2019) (CCNPs) have also been proposed as an extension of NPs, which introduces translation equivariance as an inductive bias using a convolutional extension of Deep Sets. Additionally, to enable the modeling of temporal dynamics in sequential data, a combination of RNNs and ANPs has been proposed in (Qin et al., 2019), where the hidden states of an LSTM are used to extract the feature representations. Finally, some recent extensions have also included NPs based on the Transformer architecture(Nguyen and Grover, 2022).
3.5. Hybrid Approaches

### v3

@@ 3.1.4 Probabilistic Extensions
To alleviate deterministic meta-learning’s inability to quantify uncertainty, various probabilistic approaches have been developed to improve performance and uncertainty quantification by modeling parameters as random variables and obtaining their distributions. In this direction, Bayesian meta-learning is a popular paradigm that fits expressive generative models to small datasets by incorporating inductive priors learned from related tasks. One of the first works to explore probabilistic FSL is in [47], where a hierarchical version of Variational Autoencoders (VAEs) is proposed that can be used for FSL by learning the statistics of the training data and then using them to classify new samples. The basic idea is to perform variational inference at two levels using a double variational bound where a higher-level latent variable describes datasets and a lower-level latent variable describes individual examples.
@@
Additionally, many subsequent works have focused on extending MAML to the probabilistic setting. In [69], MAML is reinterpreted as an inference for the parameters of a prior distribution in a hierarchical Bayesian model, and it is shown that the choice of update rule in the inner loop corresponds to a choice of prior over task-specific parameters. To mitigate MAML’s inaccurate point approximation, a Bayesian variant of MAML using the Laplace approximation is also proposed. In [55], MAML is extended by learning the joint posterior of initial and task-specific parameters, conditioned on each task’s support set, using structured variational inference. An amortized variational inference technique is also proposed to avoid storing separate variational distributions for each task’s initial and task-specific parameters. A Bayesian variant of MAML leveraging non-parametric variational inference has also been proposed in [241], which learns the posterior distributions of task-specific parameters using Stein Variational Gradient Descent to approximate the posterior over the weights in the final layer with a particle ensemble. However, the posterior’s expressiveness depends on the number of particles in the ensemble, leading to poor scaling regarding memory and computation costs.
@@
Apart from probabilistic MAML extensions, novel architectures have also been proposed, such as deep kernels where both feature extractor and base kernel parameters are meta-learned across tasks [162]. The deep kernel’s parameters are learned in the outer loop by maximizing the marginal likelihood across all tasks using a Gaussian process approach, obviating the need for an inner loop and leading to an efficient and straightforward closed-form solution. Finally, a novel method is introduced in [165] that meta-learns a Bayesian prior on the function space to avoid overconfident estimations. The problem is formulated as a stochastic process with different realizations, and the prior’s score function is estimated using a transformer encoder architecture.
@@
Lastly, probabilistic meta-learning can also mitigate some of metric-based meta-learning’s limitations. Specifically, metric-based meta-learning methods typically obtain a single point estimate for each class, leading to biased estimates due to a lack of data. However, a proposed extension in [249] models each class with its distribution through variational inference, allowing for the calculation of confidence and more accurate classification of novel samples. Expanding on this, in [193], computational complexity is reduced by using fewer random variables for latent class prototypes. A global model is used to learn the posterior distributions of these variables through amortized inference, and task-specific estimates of variational parameters are obtained using initial parameters learned across tasks.
3.1.5 Neural Processes
Parallel to the development of probabilistic models that extend the classic meta-learning framework, significant research effort has revolved around Neural Processes (NPs), another family of probabilistic models suitable for FSL. On a higher level, NPs aim to bridge the gap between deep neural networks (DNNs) and Gaussian Processes (GPs). DNNs excel at function approximation, yet they require retraining for each new function. On the other hand, GPs can quickly infer new functions at test time, but they are computationally expensive and sensitive to the selection of priors [64]. NPs can make accurate predictions given a few training data points and scale to complex functions and large datasets. NPs were initially developed as function estimators, suitable for modeling stochastic processes. However, their ability to model different functions using only a few samples makes them ideal for few-shot regression, where each function is a different task, and the context data constitutes the support set.
The first proposed model of this family has been Conditional Neural Processes [64] (CNPs) that uses a permutation-invariant neural network to parametrize its dependence on the observations. For each task, an aggregated mean representation of the support set is obtained and used by a decoder that models the distribution parameters of the examined target point. The model is trained end-to-end through gradient descent to maximize the conditional likelihood of a random subset of targets given a random observation set. Extending CNPs, NPs [65] introduce a more generalized approach by using a global latent variable that allows for better uncertainty modeling. Similar to CNPs, an aggregated mean representation of the support set data is obtained, which conditions the global latent variable. Finally, the decoder receives the target data and a sampled value from the conditioned global latent variable as input and predicts the output values.
Yet, despite their desired properties, these methods lack expressivity power and suffer from underfitting in complex settings. To address these limitations, Attentive Neural Processes [108] (ANPs) introduce an attention mechanism to enhance model performance. Specifically, self-attention is used to capture the interactions between support set samples, and cross-attention is used to obtain the aggregated representation. However, these attention modules also lead to increased computational complexity. Convolutional Conditional Neural Processes [68] (CCNPs) have also been proposed as an extension of NPs, which introduces translation equivariance as an inductive bias using a convolutional extension of Deep Sets. Additionally, to enable the modeling of temporal dynamics in sequential data, a combination of RNNs and ANPs has been proposed in [166], where the hidden states of an LSTM are used to extract the feature representations. Finally, some recent extensions have also included NPs based on the Transformer architecture [151].
3.2 Transfer Learning Approaches
@@
Figure 6: Two-stage model training in transfer learning. In the first stage, the model is pretrained on a large annotated dataset (base dataset), while in the second stage, it undergoes further finetuning using a smaller target dataset
@@
One of the first works to explore the efficiency of finetuning techniques for FSL has been in [28], where it is argued that finetuning methods’ performance has been underestimated in the past. Consequently, a simple baseline that utilizes a pretrained feature extractor backbone and finetunes only the classification head within each few-shot task is proposed. Classification is then performed based on the cosine similarity between the prototypes extracted from the weight matrix of the classifier head and the feature representations of the given samples. Similarly, in SimpleShot [217], a pretrained frozen backbone is also used to extract feature representations. However, a series of simple transformations such as centering and $\ell_{2}$-normalization are applied to the extracted representations, and the final classification is performed using the nearest neighbor/centroid algorithm.
@@
On a different approach, in [209], it is suggested that transferring knowledge from a base dataset to new tasks without retraining the pretrained backbone could lead to biased results due to the extraction of distorted features for out-of-distribution samples. To avoid this and prevent overfitting of the feature extractor due to the small number of samples, the whole model is finetuned using an inverted Firth Bias regularizer that regularizes the distribution of the predicted labels to be more uniform, similar to label smoothing. Full model finetuning is also proposed in [40]. In particular, the pretrained backbone is combined with a linear classifier head, and a transductive objective is introduced to finetune both the backbone and the classifier within each task, granting it the ability to adapt to different numbers of ways and shots without retraining.
@@
To address both limitations of feature distortion using a frozen pretrained backbone and the possibility of overfitting when finetuning the whole model, partial finetuning, e.g., finetuning only specific layers of the backbone network, has been proposed in [181]. To determine the finetuned layers, a genetic programming algorithm is used to learn each layer’s learning rates during finetuning, automating the whole procedure. On the other hand, in [185], FiLM modules are introduced to modulate the backbone network without retraining all of its parameters, substantially reducing the number of updated parameters. On top of it, a Naive Bayes classifier is used for prediction.
@@
Finally, various other approaches have also emerged, aiming to tackle FSL via model finetuning. In [175], the effect of invariance and equivariance is explored in the context of learning better feature representation of samples for FSL. Subsequently, the authors propose learning two additional modules alongside the classification head for enforcing equivariance and invariance via self-supervised learning, and the final model is obtained via self-distillation. Transfer learning has also been combined with meta-learning to bridge the gap between the two methods. Specifically, in [38], meta-learning is used to tune an object detection model’s hyperparameters, while finetuning is used to tackle the few-shot object detection objective in each task.
@@
Overall, finetuning methods can be a great asset in few-shot applications where simplicity and efficiency are crucial. However, there is no clear consensus on whether using a frozen backbone or finetuning it either fully or partially leads to better performance [136]. Additionally, pretraining the backbone network in a large base dataset is not always possible in some domains. Finally, cautiousness is needed when there is a domain gap between the base dataset and the novel tasks since it could hinder learning due to negative transfer.
3.3 Data Augmentation Approaches
@@
3.3.1 Basic Data Augmentation
@@
The main logic behind these approaches is to regularize the model by preventing it from overfitting to particular regions of an image and ignoring other parts. In general, this is achieved by removing specific parts of an image and replacing them with new ones, forcing the model to focus on different features. For instance, Cutout randomly removes square regions (cutouts) of the training images, while Random Erasing randomly selects and erases rectangular regions within an image and replaces them with random or predefined pixel values. On the other hand, a more recent approach, Context-Guided Augmentation [24], adds new instances of classes strategically using contextual information to simulate occlusion occurrences, with each segmented instance being selectively positioned onto the object’s bounding box in a randomized manner. Overall, image erasing methods can be seen as a form of spatial dropout that induces robustness against object occlusions and different spatial configurations. However, they suffer from the same limitations as image manipulation methods since they must be hand-picked beforehand.
3.3.2 Advanced Data Augmentation
@@
These methods aim to enhance generalization performance across diverse datasets and mitigate the risk of overfitting by integrating information from multiple sources. In Mixup [248], this is achieved by blending two or more images by taking a convex combination of their pixel values and their corresponding one-hot label encodings to create novel hybrid samples. Notably, these linear interpolations introduce diversity to the training data and assist models in learning more robust and generalized features with minimal computational overhead. CutMix [246], on the other hand, extends this concept by cutting and pasting rectangular patches from different images, fostering spatially coherent variations. More recently, Multi-stage Augmented Mixup [129] (MiaMix), a novel mixing ratio sampling method, has been proposed, consisting of four stages: random sample pairing, sampling of mixing methods and ratios, mixing mask generation and augmentation, and obtaining the mixed output. MiaMix selects a mixing method from a pool of methods and sets the mixing ratio for each mask. The masks are then augmented and mixed in the final stage. Finally, PatchMix [231] has been proposed to address the existence of biases, such as spurious correlations, in FSL. Specifically, it disentangles causal and non-causal features to improve model invariance and generalizability by augmenting original query images with images from another query while preserving the label of the additional query.
@@
Although hand-crafted DA methods have been invaluable in dealing with data scarcity and creating models invariant to various transformations, manually selecting the optimal augmentations to apply is a challenging task that often leads to suboptimal results. As a result, novel approaches have been suggested to automate this selection process by systematically finding the optimal selection policy. AutoAugment [32] formulates a search space within which each policy consists of multiple sub-policies that encapsulate different image transformations. In this setting, the most effective policy is identified through a search algorithm that maximizes validation accuracy on a target dataset based on RL principles. On the other hand, Randaugment [33] automates this selection by introducing a randomized approach. In particular, it dynamically selects a set of augmentation operations from a transformations pool and applies them with random magnitudes to each image. This approach strikes a balance between variability and consistency while it also obviates using an RL algorithm to find the optimal policy.
Figure 7: Feature data augmentation techniques extract the image’s feature representation and apply the transformation in the feature space
Figure 8: Neural style transfer methods generate synthetic images by modulating image characteristics while preserving the original semantic information
@@
While traditional DA techniques have focused on applying transformations on the raw input data, feature augmentation techniques have also been used to enhance the learning capabilities of a model by manipulating the extracted feature representations. This can involve methods such as introducing new features, transforming existing ones, or modifying specific feature values. For instance, the work in [210] proposes the Information Fusion Rectification method to identify relevant base class prototype features from the query set using cosine similarity. The method leverages the relationship between base, support, and query samples and rectifies the support set’s distribution using cosine similarity metrics as weights, generating new samples on the extracted feature space relevant to the provided task. Figure 8 shows an example of applying feature data transformations.
@@
A different way to perform DA could be by leveraging modern techniques such as Neural Style Transfer, which offers a unique approach focusing on style-based variations. In [86], the use of random style transfer is proposed by introducing randomness to texture, contrast, and color while preserving the underlying shape and semantic content of data. This is achieved by adapting an arbitrary style transfer network to perform style randomization, with target style embeddings being sampled from a multivariate normal distribution instead of being computed from a specific style image. Figure 8 illustrates an example of generating synthetic images following different styles. Finally, diffusion models have also been utilized in this context. Specifically, in [52], a pretrained CLIP encoder is used to extract latent features fed to Stable Diffusion to generate synthetic images. Data diversity is ensured by filtering generated images based on cosine similarity, while augmentation is enabled in image and feature spaces.
@@
3.4 In-Context Learning
Although not explicitly trained for it, it has been demonstrated that LLMs are capable of generalizing to novel tasks when given just a few demonstrations. This emergent few-shot ability of LLMs was first examined in [19], where scaling the model parameters and the number of training data leads to the creation of models capable of adapting to novel tasks without further parameter finetuning. Ever since, similar FSL capabilities have also been demonstrated under various model settings, such as bidirectional LLMs [163] and different tasks, including machine translation [62]. In the literature, this few-shot ability has been referred to as in-context learning [41, 258] (ICL) since the model leverages a few examples that form a demonstration context to generalize within it. In-context learning can be seen as a learning mechanism based on analogies where the model draws on its prior knowledge and forms analogies between the provided demonstrations and the query at hand.
Comparing ICL and standard FSL, one of the major differences between the two approaches is their effect on the model parameters. More precisely, the main advantage of ICL is that the model is able to adapt to a new task without having to finetune its parameters through gradient descent. Simply by structuring a prompt that provides the context for the task through input-output pairs of examples, the model can generalize to the task, while in FSL, the model’s parameters need further adaptation. As a result, ICL significantly reduces computing power and time associated with novel task adaptation. For instance, based on these few-shot capabilities, an incremental stage-wise training of language models has been proposed that enables faster and more efficient training and adaptation [173]. Moreover, another difference is that FSL models are explicitly trained to learn how to rapidly adapt to new tasks with just a few training samples, while ICL naturally emerges in LLMs without the need for explicit optimization. Finally, an additional advantage of ICL is its interpretability since the demonstrations are provided in natural language, thus facilitating the incorporation of human knowledge in them.
Undoubtedly, these desirable characteristics of ICL have rendered LLMs particularly attractive in various applications where data is limited. In [141], an autoregressive LLM is finetuned through prefix-tuning to generate new data. Consequently, the generated data is used to augment the original dataset in an attempt to render supervised learning feasible. In [35], LLMs are used as few-shot query generators to enhance information retrieval in cases where training data is absent. Notably, the proposed method manages to outperform models pretrained on large question-answering datasets by simply using eight examples as demonstration input for the LLM data generator. Furthermore, LLMs have been leveraged as few-shot abductive reasoners [182]. In that case, given a few annotated examples, their goal is to suggest possible causes for an event as part of an event sequence modeling pipeline. Finally, in recent years, ICL has also been applied to various modalities beyond text, using vision [58] and language-vision [202] models, demonstrating its potential as a new general form of FSL.
3.4.1 Theoretical Understanding of In-Context Learning
In recent years, there have been several attempts to shed light on the mechanisms that enable ICL in LLMs. On a theoretical level, various studies have analyzed how specific components within LLMs contribute to the emergence of ICL. Specifically, in [155], the authors explain how certain attention heads in the model’s architecture, called induction heads, are responsible for ICL by reproducing previously occurring token sequences for next-token prediction. More recently, this direction has been further explored using a Markov chain modeling task [46]. At the same time, another recent study has examined ICL through the lens of function vectors, which are responsible for transmitting information regarding the demonstration examples across the transformer modules [200].
Given the similarities between ICL and meta-learning for FSL, in terms of their ability to adapt to novel tasks using a minimal number of samples, various studies have attempted to frame ICL as a type of gradient-based meta-learning, where the forward pass of the transformer modules in ICL approximates performing finetuning within each task’s support set in FSL [206]. Specifically, in [34], GPT-based models are viewed as meta-optimizers and ICL as implicit model finetuning based on a recasting of transformer attention as a meta-optimization module. However, recent studies have suggested that these similarities may not be sufficient to fully explain ICL due to differences between the inner workings of ICL and gradient descent [39], necessitating further study in that area.
Another line of theoretical research on ICL has focused on analyzing it as a form of function regression learning, where the model learns how to fit a regression function to the provided demonstration examples. In that direction, most existing works have focused on learning linear regression functions from the given examples [63, 3]. At the same time, a more recent study has also extended this framework to a broader setting by introducing a compositional structure that combines learning a fixed representation function with a linear regression function [70].
Lastly, in [230], ICL is presented as a form of implicit Bayesian inference, and the effect of the training data distribution is underlined. The same framing has also been combined with the regression function learning [158] and gradient-based meta-learning [89] frameworks in an attempt to provide a broader and unified perspective of how ICL works.
3.4.2 Empirical Insights on In-Context Learning
Parallel to the works providing a theoretical framework for ICL, various studies have focused on identifying and quantifying the effect of different influencing factors from an empirical perspective. More specifically, model size [101] has been pinpointed as an influencing factor, yet it is inadequate to illustrate the intrinsic properties that lead to the emergence of ICL. Similarly, other model properties dubbed as possible influencing factors have been computation [222], and the objectives selected during model pretraining [198]. In terms of the data used during model pretraining, attributes such as their domain [183], distribution [22], and diversity [171] have been examined, underscoring the importance of the pretraining data for the emergence of ICL.
Since LLMs are not explicitly trained to perform ICL, this emergent property can also be influenced by various factors during inference. Specifically, attributes of the examples used as demonstrations in the ICL setting, such as their diversity [6], their similarity with the test query [133], and their ordering [134] have been shown to impact ICL performance. Moreover, another critical factor during inference affecting ICL is the accuracy of the mapping between inputs and labels in the demonstrations and the samples used during pretraining [112].
However, there are cases where findings seem to be contradictory regarding the actual impact of the various examined factors. For instance, while in [98] a correlation between the relevance of the pretraining data and the examined downstream task is identified, in [72], the authors show that this factor does not influence the effectiveness of ICL. Additionally, although initial findings have demonstrated that the input-label mapping does not influence ICL [145], subsequent studies have revealed that a consistent input-label mapping between pretraining data and the demonstrations used in ICL can lead to improved performance [112].
3.4.3 Boosting In-Context Learning with Meta-Learning
An interesting aspect of ICL is that despite being itself a form of FSL, it has recently been combined with other FSL paradigms and, most notably, meta-learning to enhance its performance. Specifically, ICL is related to prompt learning, which allows for the use of learnable prompts instead of hand-crafted ones. In this context, meta-learning can be utilized to efficiently learn better prompts across various downstream tasks, thus enhancing the performance of ICL. Although LLMs already demonstrate implicit few-shot capabilities, their performance in FSL scenarios can be further increased through explicit training on few-shot formulated tasks. For instance, in [144], a meta-training scheme based on formulating few-shot tasks is proposed to enhance the ICL capabilities of LLMs. Additionally, in [79], an approach adapting MAML for prompt tuning is proposed, which meta-learns an effective prompt initialization that enables fast adaptation to multiple downstream tasks. Lastly, a novel approach has recently been proposed aiming to improve ICL through meta-ICL [31]. More specifically, although ICL allows for rapid adaptation to downstream tasks using a minimal number of demonstrations, different tasks may benefit from different priors and ICL strategies. By introducing an additional level where ICL is performed across sets of downstream tasks, the model is able to adapt its priors and learning strategies and perform ICL more effectively within each group of tasks.
3.4.4 Limitations
Although the benefits of ICL have been both theoretically explored and practically demonstrated, there are still various limitations pertinent to its inherent nature. As discussed earlier, ICL can be unstable, influenced by factors such as the prompting format, the examples provided, and even their ordering [254]. To address these issues, different calibration methods have been proposed to make LLMs less sensitive to these parameters. For instance, in [254], various types of biases in prompts are identified, and a contextual calibration approach is proposed based on content-free test inputs. In [73], prototypical networks are utilized to calibrate the decision boundary of LLMs for classification in the ICL setting. This is achieved by creating prototypical clusters based on the support set samples using a Gaussian Mixture Model. Consequently, when combined with this type of calibration, LLMs demonstrate increased robustness regarding prompt ordering, class imbalance, and size of estimate set. Lastly, a recent study has posited that ICL can be viewed as a byproduct of using nonlinear or discontinuous metrics rather than an emergent property of LLMs [178]. Overall, this lack of clear consensus regarding a comprehensive theoretical and empirical justification of ICL highlights the need for further research to shed light on this novel and distinct form of FSL.
4 Few-Shot Learning Beyond Supervised Learning Settings